# A Search for Classical Subsystems in Quantum Worlds

### v2

@@ IV.4 Analysis
and pick, for simplicity, two particular but arbitrary indices $p$, $q$ such that $[H_{s}^{p},H_{s}^{q}]=0$ Then, any initial state of the form $|\psi\rangle=|s_{i}\rangle(\alpha|p\rangle_{e}+\beta|q\rangle_{e})$ (where $|s_{i}\rangle$ is a mutual eigenstate of $H_{s}^{p/q}$) will remain unentangled. Of course, while we picked just two commuting $H_{s}^{(i)}$ terms for simplicity, the argument is quite general; all that is needed is for the environment state to project the total state into a subspace that admits a pointer state solution.
The form of the Hamiltonian in Eq. (20) also makes it clear that, for a fixed system size, the larger the environment the more such solutions will exist. This is because there are $d_{e}$ terms and if each of the $H_{s}^{(i)}$ operators is generated randomly the likelihood that any two operators will commute (or approximately commute) increases. As shown in Fig. 3 d., clearly the trained state remains unentangled throughout the evolution while the other states, including the state made from randomizing the environment state, generally do not. Although in this figure the ‘furnace’ solution appears to resemble the ‘approximate eigenstate’ solution, the distinguishing feature can be seen in Fig. 5 which shows that the purity of the reduced density matrix of the trained state is robust to arbitrary changes within a particular subspace.
Intuitively, this solution can be understood in the following physical sense: consider a pendulum in the lab that undergoes constant environment monitoring and thus stays in the pointer state as it oscillates (as is expected of a subsystem). But now the lab assistant turns up the temperature of the room, as in a furnace, so that the pendulum incinerates and its atoms are dispersed amongst the various degrees of freedom available in the room. Here, though the Hamiltonian has remained the same, changing the environment state has drastically altered the description of the system to one that no longer has a pointer state.
@@
These solutions sustain a moderately high purity for $1-2$ orders of magnitude longer than the typical decoherence time. From the algorithm’s perspective such solutions can be thought of as a result of poor training or poor generalizability. In the former case, the algorithm finds a combination of initial state and factorization such that the solution only roughly approximates one of the other “perfect” solutions above. However, the convergence to such a solution is slow and the algorithm hits a time-wall before it can get sufficiently close to such a solution. Poor generalizability occurs when the algorithm is able to sufficiently minimize the cost for the training steps, but that the solution does not generalize well beyond those finite times. In this case, the situation resembles the purity plots shown in the left panel of Fig. 1. As mentioned in Section III, since we take our maximum training time to be $\approx 10^{2}t_{\rm dec}$, it makes sense that the failure to generalize manifests beyond this domain.
@@
Such ‘extended coherence time’ solutions also occur in the case of fixed input state optimization (i.e. $A=\mathbb{1}$). This is because, in fixing $A$, the algorithm loses control over the degrees of freedom required to sufficiently minimize the cost and to ensure that the purity of the reduced density matrix remains high for times much greater than $t_{\rm dec}$. We find these solutions for all classes of Hamiltonians and they may be interpreted as short-lived subsystems. That is, as a set of dynamics which, a priori, appears highly entangling but can be viewed in another factorization as temporarily pertaining to subsystems. Ultimately, we lack an analytic understanding of these solutions and caution from reading too much into the numerics shown here. Firstly, the purities obtained in this case are substantially lower than in the case where we optimise over the initial state (see the right panel of Fig. 8). Secondly, it is unclear how these solutions generalize to larger systems. Thus, rather than highlighting new physics, perhaps the fleeting subsystems found here are akin to shapes found in the sky when cloud watching. On the other hand, these solutions might most closely reflect the fate of all subsystems in a dark energy dominated universe.
Figure 5: In the furnace Hamiltonian, a pointer state for the system exists only if the environment is in a particular subspace of $\mathcal{H}_{e}$ (unlike the Block diagonal case, c.f. Fig. 3b, where the pointer state is indifferent to the choice of $|\phi\rangle_{e}$). Here we depict this phenomenon by perturbing the trained state $|\psi\rangle_{w}$ (turquoise) such that the new state $|\psi^{\prime}\rangle_{w}$ (purple) has support from the same energy eigenvectors as those that contribute to the trained state (top panel). The bottom panel verifies that such a state indeed remains a product state at late times. Note that $|\psi\rangle$ and $|\psi^{\prime}\rangle$ share the same system state in $\mathcal{H}_{s}$ (up to an overall phase). We show, for comparison, another state where $|\phi\rangle_{e}$ is randomized leading to a large increase in the linear entropy in the typical decoherence time (the furnace “incinerates” in the analogy of the text in Section IV.4(d)).
@@ V Many more worlds
Everett’s relative state interpretation of quantum mechanics [43, 44, 45] was meant to address the system-observer dichotomy, where the observer/apparatus is treated as a classical object (and many textbooks continue to present this dichotomous exposition). The relative state interpretation, and its various extensions in the many-minds, de Witt’s many-worlds [46], or Zurek’s existential [19] interpretations, posits a global wave function that evolves unitarily and definite perceived outcomes result from the state of the observer becoming correlated with the state of the system. It is in this sense that the relative state interpretation may be thought of as a literal reading of unitary quantum mechanics. For concreteness, consider a system in the state
@@
where $|a_{r}\rangle$ is the “ready” state of the apparatus. The relative state interpretation posits that no definite state can be ascribed to the system or the apparatus but, instead, the state of the apparatus (or the observer), say $|a_{1}\rangle$, should be understood as being correlated with the state $|1\rangle$ of the system. Then, each of the two terms of Eq. (22) correspond to the two “branches” of the wave-function. But why should the branches be described by the particular choice of basis used in Eq. (22)? Consider instead the system and apparatus bases,
|±$|\pm\rangle_{s}=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)\textrm{ and }|a_{\pm}\rangle=\frac{1}{\sqrt{2}}(|a_{0}\rangle\pm|a_{1}\rangle);,$
so that the final global state in Eq 22 takes the form,
@@
This ambiguity in the choice of basis between Eq. (22) and Eq. (23) poses the questions: what exactly has the apparatus “measured”? In the language of the many worlds interpretation, which basis (that of Eq. (22) or 23) describes the post-measurement world? [^9] The fact that the two terms in Eq. (21) have equal coefficients is a choice that simplifies our mathematical expressions, but there are equivalent issues for the general case. Einselection resolves this preferred basis problem [12] by considering the interaction of the apparatus with the environment which, depending on the details of the Hamiltonian, singles out a set of preferred basis so that it is, say, the $\{a_{0},a_{1}\}$ basis that remains robust to entanglement with the environment and that the state of the system should be inferred relative to this basis [14].
@@
Thus, einselection provides a way to dynamically select the preferred basis of the subsystems. However, the presence of pointer states is crucial and requires specially constructed Hamiltonians (whether through the predictability sieve strategy [25] or by virtue of vanishing commutators in the self and interaction Hamiltonians [14]). In particular, it requires one to specify the individual subsystems that the Hamiltonian describes.
@@
In this paper we set out to answer the question if the division into subsystems can be derived, as opposed to assumed, from the spectrum of the Hamiltonian [^10] This is similar in spirit to D. Deutsch’s “interpretation basis” approach [8] where he proposes a resolution to the preferred basis problem that can, in turn, be used to address the issue of factorization into subsystems (see Eq. 50 and surrounding discussion therein). However, see Ref. [47] for crucial arguments against the interpretational basis approach. . In Section IV.4, we have shown that there always exists at least one such factorization – that in which the global eigenbasis is separable, that admits pointer states and hence can exhibit einselection (in the sense of Eq. 22 above) and that for “real-world” Hamiltonians, several other choices also exist. But this introduces a new issue: the block diagonal factorization is available in addition to the other factorizations that also support einselection (such as the coherent state wavepackets of the ACL model discussed in Section II). What then favors one factorization over the other? We stumble upon a “global” analog of the preferred basis problem: if multiple TPS admit decoherent histories why should the world branch in one factorization over the other? Why should the physicist’s choice of global basis in writing down the state determine the branches of the worlds? Moreover, the resulting ambiguity is in a sense more unsettling than the preferred basis ambiguity of the system subspace. The various choices of TPS describe many “realms”: the set of worlds described by the fiducial factorization have a different ontology than the set of worlds in the realm described by the destination factorization. The harmonic oscillator of the ACL is far from an SHO in a different system-environment split. And while both realms share the same global energy spectrum, the scene described by a central spin surrounded by environment qubits in an external magnetic field is radically different than the fermionic degrees of freedom with hopping terms obtained by taking a Jordan-Wigner transform. Note also that all these realms exist at fixed system and environment size, which has been the focus of this work. Allowing additional freedom in the sizes of the individual Hilbert space factors, which in principle is unobjectionable for the global state since two Hilbert spaces of the same dimensions are isomorphic, will only add to the diversity of the realms. Furthermore, while we have focused on only bipartite factorizations, the case of multipartite subsystems may further add to this diversity, especially if one considers $\dim{(\mathcal{H}_{w})}$ to be highly factorizable as appears to be the case for our Universe [48].
@@
This ontological ambiguity between the realms is unsettling, and some might want to refer to a particular “physical” factorization as describing “real” subsystems, and the other factorizations describing subsystems which are “virtual” or in some sense less “real”, we reserve judgment on the “reality” of eccentric realms  [^11] Coleman poses the same issue in the context of bosonization [34], “One of the most striking features of the results established here is that a theory which is ‘obviously’ a theory of fermions is equivalent to a theory of bosons… That these theories are equivalent in any sense seems too preposterous to believe…” . One might argue that the “physical” factorizations have some notion of locality either in the pointer state as in [2, 5], at the level of the Hamiltonian [4], or in the set of observables [1]. While there is some merit to this argument, we remain curious if the “local” factorizations are in some deeper sense preferred over the other ones and whether one can derive this preference from something fundamental rather than inflicting this classical prejudice onto the Hilbert space. Assuming locality is principally not much different than assuming a particular factorization of the Hamiltonian as is done implicitly in most formulations, and explicitly in Wallace’s three postulates [49, Chapter 1], of QM.
@@
Similar prejudices enter into arguments that try to minimize $H_{\rm int}$ in an effort to look for a preferred factorization [3, 7]. But there is nothing in the einselection formalism (or the relative state interpretations) that presupposes a TPS that minimizes the interaction Hamiltonian. [^12] A general discussion of einselection behaviors under different interaction strengths appears in [32] and several cases are illustrated in concrete terms for the ACL model in [20]. Attempts to search for such a preferred TPS by minimizing the interaction strength are misguided because important physical cases are described by a dominant $H_{\rm int}$ term and in those cases einselection occurs not inspite of but because of the strong interaction (see Quantum Measurement Limit in Section IV.2).
@@
While others have put forth “in-principle” algorithms to search for quasi-classical factorizations [5, 8], we have been successful in practically implementing an algorithm to explore the many realms that coexist in worlds with a given global energy spectrum. What remains missing is some measure on the space of these realms serving much the same purpose as that served by the predictability sieve for the preferred basis problem, or by the Born rule for the many world branches (to suppress the “maverick” worlds [50]). As we explain in Appendix B, one such measure is induced by our algorithm based on the minimization of the cost Eq. (10). However it is unlikely that such a measure is of any physical significance.
The abundance of coexisting realms, and the absence of a physical measure to discriminate among them appears to be related to issues that have been explored in the context of the consistent histories formalism, where alternative sets of complete projectors lead to different coexisting narratives [51, 52, 53, 54] (“alternate facts”, in the language of [55]). However, those discussions typically do not explore as wide a range of tensor product structures as we consider in this paper. The need for more information than is contained in a Hamiltonian to determine a unique tensor product structure has also been investigated from a mathematical perspective in [9, 56, 57]. Finally, we note that there are also some parallels with discussions of the “clock ambiguity” in the context of quantum gravity [58].

### v3

@@ IV.3 Analysis
and pick, for simplicity, two particular but arbitrary indices $p$, $q$ such that $H_{s}^{p},H_{s}^{q}]=0$ . Then, any initial state of the form $|\psi\rangle=|s_{i}\rangle(\alpha|p\rangle_{e}+\beta|q\rangle_{e})$ (where $|s_{i}\rangle$ is a mutual eigenstate of $H_{s}^{p}$ and $H_{s}^{q}$) will remain unentangled. Of course, while we picked just two commuting $H_{s}^{(i)}$ terms for simplicity, the argument is quite general; all that is needed is for the environment state to project the total state into a subspace that admits a pointer state solution.
Figure 4: In the furnace Hamiltonian, a pointer state for the system exists only if the environment is in a particular subspace of $\mathcal{H}_{e}$ (unlike the Block diagonal case, c.f. Fig. 2b, where the pointer state is indifferent to the choice of $|\phi\rangle_{e}$). Here we depict this phenomenon by perturbing the trained state $|\psi\rangle_{w}$ (turquoise) such that the new state $|\psi^{\prime}\rangle_{w}$ (purple) has support from the same energy eigenvectors as those that contribute to the trained state (top panel). The bottom panel verifies that such a state indeed remains a product state at late times. Note that $|\psi\rangle$ and $|\psi^{\prime}\rangle$ share the same system state in $\mathcal{H}_{s}$ (up to an overall phase). We show, for comparison, another state where $|\phi\rangle_{e}$ is randomized leading to a large increase in the linear entropy in the typical decoherence time (the furnace “incinerates” in the analogy of the text in Section IV.3(d)).
The form of the Hamiltonian in Eq. (21) also makes it clear that, for a fixed system size, the larger the environment the more such solutions will exist. This is because there are $d_{e}$ terms and, if each of the $H_{s}^{(i)}$ operators is generated randomly, the likelihood that any two operators will commute (or approximately commute) increases. As shown in Fig. 2d, clearly the trained state remains unentangled throughout the evolution while the other states, including the state made from randomizing the environment state, generally do not. Although, in this figure, the ‘furnace’ solution appears to resemble the ‘approximate eigenstate’ solution, the distinguishing feature can be seen in Fig. 4 which shows that the purity of the reduced density matrix of the trained state is robust to arbitrary changes within a particular subspace spanned by a subset of the energy eigenstates.
Intuitively, this solution can be understood in the following physical sense: consider a pendulum in the lab that undergoes constant environment monitoring and thus stays in the pointer state as it oscillates (as is expected of a subsystem). But now the lab assistant turns up the temperature of the room, as in a furnace, so that the pendulum incinerates and its atoms are dispersed amongst the various degrees of freedom available in the room. Here, though the Hamiltonian has remained the same, changing the environment state has drastically altered the description of the system to one that no longer has a well-defined classical subsystem.
@@
These solutions sustain a moderately high purity for 121-2 orders of magnitude longer than the typical decoherence time. From the algorithm’s perspective such solutions can be thought of as a result of poor training or poor generalizability. In the former case, the algorithm finds a combination of initial state and factorization such that the solution only roughly approximates one of the other “perfect” solutions above. However, the convergence to such a solution is slow and the algorithm hits a time-wall before it can get sufficiently close to such a solution. Poor generalizability occurs when the algorithm is able to sufficiently minimize the cost for the training steps, but that the solution does not generalize well beyond those finite times. In this case, the situation resembles the purity plots shown in the left panel of Fig. 6. As mentioned in Section III, since we take our maximum training time to be $\approx 10^{2}t_{\rm dec}$, it makes sense that the failure to generalize manifests beyond this domain.
@@
Such ‘extended coherence time’ solutions also occur in the case of fixed input state optimization (i.e. $A=\mathbb{1}$). This is because, in fixing $A$, the algorithm loses control over the degrees of freedom required to sufficiently minimize the cost and to ensure that the purity of the reduced density matrix remains high for times much greater than $t_{\rm dec}$. We find these solutions for all classes of Hamiltonian and they may be interpreted as short-lived subsystems. That is, as a set of dynamics which, a priori, appears highly entangling but can be viewed in another factorization as temporarily pertaining to subsystems. Ultimately, we lack an analytic understanding of these solutions and caution from reading too much into the numerics shown here. Firstly, the purities obtained in this case are substantially lower than in the case where we optimise over the initial state (see the right panel of Fig. 8). Secondly, it is unclear how these solutions generalize to larger systems. On the other hand, these solutions might most closely reflect the fate of all subsystems in a dark energy dominated universe, which could ultimately end up as a thermalized de Sitter space.
To summarize, we have found five different categories of solution to our optimization problem, each with its distinct set of associated behaviors. The block diagonal destination Hamiltonian is always available as a TPS that supports einselection, even for random Hamiltonians. Thus, one need not construct special Hamiltonians to demonstrate decoherence. The decoupled case can be realized for certain structured Hamiltonians. The other three cases are intriguing in their own right and each offers its own set of open questions (such as how important is it to have a complete pointer basis, and what patterns of energy eigenstate entropies are conducive to classical dynamics). Perhaps most intriguing are cases where these different destination Hamiltonians can be found for the same set of eigenvalues.The implication is that different types of subsystem coexist, at least formally, within the same Hamiltonian. We next explore some implications of this result for quantum foundations.
@@ V Many more worlds
Our results appear to have provocative implications for certain interpretations of quantum mechanics. We will not make value judgments on the interpretations themselves but rather point out that the presence of different TPS’s that coexist, describing alternative semiclassical subsystems of the same larger system leads to an ambiguity in the global basis with respect to which the branches of the wavefunction may be identified with classical reality.
Everett’s relative state interpretation of quantum mechanics [everett1956Thesis, everett1957relative, wheeler1957assessment] was meant to address the system-observer dichotomy, where the observer/apparatus is treated as a classical object (and many textbooks continue to present this dichotomous exposition). The relative state interpretation, and its various extensions in de Witt’s many-worlds interpretation [dewitt1973many-worlds] or Zurek’s existential [zurek1998decoherence-existential-interpretation-rough-guide] interpretation, posits a global wave function that evolves unitarily and definite perceived outcomes result from the state of the observer becoming correlated with the state of the system. It is in this sense that the relative state interpretation may be thought of as a literal reading of unitary quantum mechanics. For concreteness, consider a system in the state
@@
where $|a_{r}\rangle$ is the “ready” state of the apparatus. The relative state interpretation posits that no definite state can be ascribed to the system or the apparatus but, instead, the state of the apparatus (or the observer), say $|a_{1}\rangle$, should be understood as being correlated with the state $|1\rangle$ of the system. Then, each of the two terms of Eq. (23) correspond to the two “branches” of the wave-function.
However, the relative interpretation then faces the so-called ‘preferred basis’ problem. Namely, why should the branches be described by the particular choice of basis used in Eq. (23)? Consider instead the system and apparatus bases,
$|\pm\rangle_{s}=\frac{1}{\sqrt{2}}(|0\rangle\pm|1\rangle)\textrm{ and }|a_{\pm}\rangle=\frac{1}{\sqrt{2}}(|a_{0}\rangle\pm|a_{1}\rangle);,$
so that the final global state in Eq. (23) takes the form,
@@
This ambiguity in the choice of basis between Eq. (23) and Eq. (24) poses the questions: what exactly has the apparatus “measured”? In the language of the many worlds interpretation, which basis (that of Eq. (23) or (24)) describes the post-measurement world? [^11] The fact that the two terms in Eq. (22) have equal coefficients is a choice that simplifies our mathematical expressions, but there are equivalent issues for the general case.
Einselection resolves this preferred basis problem [zurek2003decoherence-and-the-quantum-origins] by considering the interaction of the apparatus with the environment. Depending on the details of the Hamiltonian, this apparatus-environment action singles out a set of preferred basis so that, say, the $\{a_{0},a_{1}\}$ basis are pointer states that remains robust to entanglement with the environment. The state of the system should be inferred relative to this basis. In this manner, einselection provides a way to dynamically select the ontologically significant branches of the wavefunction for relative state interpretations.
@@
However, the presence of pointer states is a crucial assumption in previous work on einselection and requires specially constructed Hamiltonians (whether through the predictability sieve strategy [zurek1993preferred-predictability-sieve] or by virtue of vanishing commutators in the self and interaction Hamiltonians [zurek1981pointer]). In particular, it requires one to specify the individual subsystems that the Hamiltonian describes.
@@
In this paper, we set out to answer the question of whether the division into subsystems can be derived, as opposed to assumed, from the spectrum of the Hamiltonian [^12] This is similar in spirit to D. Deutsch’s “interpretation basis” approach [deutsch1985quantum-interpretation-basis] where he proposes a resolution to the preferred basis problem that can, in turn, be used to address the issue of factorization into subsystems (see Eq. 50 and surrounding discussion therein). However, see Ref. [foster1988recent-interpretation-basis] for crucial arguments against the interpretational basis approach. . In Section IV.3, we have shown that there always exists at least one such factorization – that in which the global eigenbasis is separable, that admits pointer states and hence can exhibit einselection (in the sense of Eq. 23 above). Moreover, for “real-world” Hamiltonians, several other choices also exist. But, this introduces a new issue: the block diagonal factorization is available in addition to the other factorizations that also support einselection (such as the coherent state wavepackets of the ACL model discussed in Section II). What then favors one factorization over the other?
Our findings in this paper thus highlight a “global” analog of the preferred basis problem: if multiple TPSs admit consistent histories why should the world branch in one factorization over the other? Why should the physicist’s choice of global basis in writing down the state determine the branches of the worlds? Moreover, the resulting ambiguity is in a sense more unsettling than the preferred basis ambiguity of the system subspace. The various choices of TPS describe many “realms”: the set of worlds described by the fiducial factorization have a different ontology than the set of worlds in the realm described by the destination factorization. The harmonic oscillator of the ACL is far from an SHO in a different system-environment split. And while both realms share the same global energy spectrum, the scene described by a central spin surrounded by environment qubits in an external magnetic field is radically different than the fermionic degrees of freedom with hopping terms obtained by taking a Jordan-Wigner transform.
While in this paper we have focused on the case of bipartite factorizations, this new preferred basis problem (in contrast to the resolution of the original preferred basis problem) looks unlikely to be resolved by expanding to tripartite/multipartite splits. Namely, the case of multipartite subsystems may further add to this diversity of the realms, especially if one considers $\dim{(\mathcal{H}_{w})}$ to be highly factorizable as appears to be the case for our Universe [eakins2002factorization-microsingularity].
@@
Thus this ontological ambiguity between the realms is unsettling, and some might want to refer to a particular “physical” factorization as describing “real” subsystems, and the other factorizations describing subsystems which are “virtual” or in some sense less “real”, we reserve judgment on the “reality” of eccentric realms  [^13] Coleman poses the same issue in the context of bosonization [coleman1975quantum-bosonization], “One of the most striking features of the results established here is that a theory which is ‘obviously’ a theory of fermions is equivalent to a theory of bosons… That these theories are equivalent in any sense seems too preposterous to believe…” . One might argue that the “physical” factorizations have some notion of locality either in the pointer state as in [piazza2010glimmers, carroll2021quantum-singh], at the level of the Hamiltonian [Cotler:2017abq], or in the set of observables [zanardi2004quantum-lloyd-lidar]. While there is some merit to this argument, we remain curious if the “local” factorizations are in some deeper sense preferred over the other ones and whether one can derive this preference from something fundamental rather than inflicting this classical prejudice onto the Hilbert space. Assuming locality is principally not much different than assuming a particular factorization of the Hamiltonian as is done implicitly in most formulations, and explicitly in Wallace’s three postulates [wallace2012emergent-multiverse, Chapter 1] of quantum mechanics.
@@
Similar prejudices enter into arguments that try to minimize $H_{\rm int}$ in an effort to look for a preferred factorization [tegmark2015consciousness, zanardi2022operational]. But there is nothing in the einselection formalism (or the relative state interpretations) that presupposes a TPS that minimizes the interaction Hamiltonian. [^14] A general discussion of einselection behaviors under different interaction strengths appears in [schlosshauer2007decoherence] and several cases are illustrated in concrete terms for the ACL model in [albrecht2023adaptedcalmodel]. Attempts to search for such a preferred TPS by minimizing the interaction strength are misguided because important physical cases are described by a dominant $H_{\rm int}$ term and in those cases einselection occurs not in spite of but because of the strong interaction (see Quantum Measurement Limit in Section IV.1).
@@
While others have put forth “in-principle” algorithms to search for quasi-classical factorizations [deutsch1985quantum-interpretation-basis, carroll2021quantum-singh], we have been successful in practically implementing an algorithm to explore the many realms that coexist in worlds with a given global energy spectrum. What remains missing is some measure on the space of these realms serving much the same purpose as that served by the predictability sieve for the preferred basis problem, or by the Born rule for the many world branches (to suppress the “maverick” worlds [dewitt1970quantum-Physics-Today]). As we explain in Appendix D, one such measure is induced by our algorithm based on the minimization of the cost Eq. (12). However, it is unlikely that such a measure is of any physical significance.
@@
Let us end this section by commenting on potential links between our findings here and investigations elsewhere. Firstly, the abundance of coexisting realms, and the absence of a physical measure to discriminate among them appears to be related to issues that have been explored in the context of the consistent histories formalism, where alternative sets of complete projectors lead to different coexisting narratives [Albrecht:1992rs, Albrecht:1992uc, Dowker:1994ac, Brun_Hartle_1999, wallace2002worlds-consistent-histories] (“alternate facts”, in the language of [Albrecht:2021sic]). However, those discussions typically do not explore as wide a range of tensor product structures as we consider in this paper. The need for more information than is contained in a Hamiltonian to determine a unique tensor product structure has also been investigated from a mathematical perspective in [Stoica:2021bub, Stoica:2021rqi, Stoica:2024nhh]. Finally, we note that there are also some parallels with discussions of the “clock ambiguity” in the context of quantum gravity [Albrecht:2007mm].